OpenStudy (anonymous):

Please help me to determine d/dx(4√x-8/√x+πx^3)

OpenStudy (anonymous):

The best way to do such problems is to do a substitution (at least when you're starting) u = 4√x-8 v = √x+πx^3 What is (u/v)' ?

OpenStudy (anonymous):

Hint: (1/v)' = -1/v^2 Now use the product rule Or look up the quotient rule.

OpenStudy (anonymous):

telliott99 can you go further please?

OpenStudy (anonymous):

Find \[\frac{d}{dx} (\frac{4 \sqrt{x} - 8}{\sqrt{x} + \pi x^3})\] Calculate the derivatives of numerator and denominator \[\frac{d}{dx} (4 \sqrt{x} - 8) = \frac{2}{\sqrt x}\] \[\frac{d}{dx} ({\sqrt{x} + \pi x^3}) = \frac{1}{2 \sqrt x} + 3 \pi x^2\] Use the quotient rule \[(u/v)' = (v u' - u v')/v^2\] Substituting, the numerator is: \[(\sqrt{x} + \pi x^3)(\frac{2}{\sqrt x}) - (4 \sqrt {x} - 8) (\frac {1}{2 \sqrt(x)} + 3 \pi x^2)\] And the denominator is: \[(\sqrt {x} + \pi x^3)^2\] After that it's just algebra. Ugly, but algebra.